OpenStudy (anonymous):

The graph of g(x) is f(x) translated to the left 8 units and up 2 units. What is the function rule for g(x) given f(x) = x²?

OpenStudy (ranga):

When you move a graph to the right by "a" units, the x on the right hand side of the function should be replaced with (x-a) If you move a graph up by "b" units, you should add "b" to the right hand side of the function.

OpenStudy (ranga):

Moving to the left or down follows similar pattern: When you move a graph to the Left by "a" units, the x on the right hand side of the function should be replaced with (x+a) If you move a graph down by "b" units, you should subtract "b" from the right hand side of the function.

OpenStudy (anonymous):

(x-+8)² + 2 (x-8²) + 2 So like that?

OpenStudy (ranga):

g(x) = (x - (-8))^2 + 2 g(x) = (x + 8)^2 + 2

OpenStudy (anonymous):

Is that there is all to it? that was easy ^-^

OpenStudy (ranga):

yes, that is all they are asking for here.

OpenStudy (anonymous):

Thanks for helping me!

OpenStudy (ranga):

you are welcome.